OpenStudy (anonymous):

How do you figure this out? 4(c+5)-f^4 c=-1, f=4

OpenStudy (zzr0ck3r):

4(-1+5)-4^4 = 4(4)-256 = 16-256 = -240

OpenStudy (phi):

These just mean: replace the letters with the numbers they give you. Then do the arithmetic